# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ VII Conclusions
As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) ‘and’-definitions, (ii) ‘add’-definitions, and (iii) ‘or’-definitions. In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant. Whereas, in the case of add-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants. Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants. The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014) which, in the case of bimatrix games, is equivalent to another concept called Nash–Pareto pair Hofbauer and Sigmund (1998).
@@
If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games. One finds that ESS’s of and- and add- definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed. However, when it comes to the desirable property that ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, 2ESS need not possess it. Using linear stability analysis and the information-theoretic concept of KL-divergence as a Lyapunov function, we have shown that in this respect and-definitions are most complying. When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed. While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure 2ESS need not be so unlike ESS’s of and- and add-definitions. Here, it must be pointed out that the rigorous proof that mixed 2ESS in asymmetric hypermatrix games is locally asymptotically fixed point is merely a conjecture seeking a proof that we failed to provide in this paper. Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
The most important take-home message, we would like to convey is that the and-definitions are the most relevant and robust as far as the issue of attainment of ESS is concerned. The significance of 2ESS is solely in asymmetric games, where it corresponds to a locally asymptotically stable interior fixed point, since some 2ESS’s may not be ESS of and- and add-definitions.
@@
We conclude by highlighting potential avenues for extending the findings of this paper. While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question. Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al. (2022) in non-matrix population games remains unexplored. Exploring the realm of asymmetric games in finite Nowak et al. (2004); Sekiguchi and Ohtsuki (2015) populations to locate an appropriate definition of ESS is also an interesting problem. Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al. (2022), opens up another avenue for investigation.
@@ Acknowledgements.
Data availability statement Data sharing is not applicable to this article as no new data is created or analyzed in this study.
                                                                 Declarations
Conflict of interest The authors declare that they have no conflict of interest.
@@
A subsequent sentence in the scheme of proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second. Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second. This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper. For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed Nash equilibrium), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$. Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$. To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if first component is zero then automatically second component is also zero. Let us see explicitly why this is so.
We know that any $2\times 2$ bimatrix game can be expressed in the form of either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V). For a rescaled zero sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998). This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
The theorem, as is present, in Ref. Cressman (1996) erroneously establishes the asymptotic stability of mixed 2ESS for even bimatrix games. We can assert that this oversight in Ref. Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
@@ Appendix B ESS under Definition 2e for bihypermatrix games can only be pure
Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$. Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\epsilon}_{x},\mathbf{\hat{y}}+\bm{\epsilon}_{y})\equiv((\hat{x}+\epsilon_{x},1-\hat{x}-\epsilon_{x}),(\hat{y}+\epsilon_{y},1-\hat{y}-\epsilon_{y}))$. We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\epsilon_{x}$ and $\epsilon_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative. Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\epsilon}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\epsilon}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (83)
@@
$=$ $-\epsilon_{x}\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)-\epsilon_{y}\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
$=$ $-\epsilon_{x}\left(b_{1}\epsilon_{y}+b_{2}\epsilon_{y}^{2}+\cdots+b_{\gamma_{x}}\epsilon_{y}^{\gamma_{x}}\right)-\epsilon_{y}\left(c_{1}\epsilon_{x}+c_{2}\epsilon_{x}^{2}+\cdots+c_{\nu_{x}}\epsilon_{x}^{\nu_{y}}\right).$
@@
Now, to the leading order, the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\epsilon_{x}\epsilon_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\epsilon_{x}$ and $\epsilon_{y}$. For similar reason, in case $-(b_{1}+c_{1})\epsilon_{x}\epsilon_{y}=0$, the next order term in $\epsilon_{x}$ and $\epsilon_{y}$ of the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) leads to same conclusion, and so on. In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games. We remark the line of proof adopted herein can be extended to the case of mixed ESS under Definition 1e.

### v2

@@ VII Conclusions
As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) and-definitions, (ii) sum-definitions, and (iii) or-definitions. In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant. Whereas, in the case of sum-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants. Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants. The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014) which, in the case of bimatrix games, is equivalent to another concept called Nash–Pareto pair Hofbauer and Sigmund (1998).
@@
If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games. One finds that ESS’s of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed. However, when it comes to the desirable property that an ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is non-strict NE. When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed. While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points. Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
The most important take-home message, we would like to convey is that while strict NE is ESS under all three classes of definitions, other ESS’s need not simultaneously satisfy all the classes. Not only that, the latter ESS’s may not always be locally asymptotically stable under replicator equation. In a given situation, it may be practical to adopt that definition which renders ESS locally asymptotically stable; after all, if an ESS is not reached dynamically starting from a mutant-invaded state, then it doesn’t seem to serve its purpose. The asymptotic approach towards ESS state is also compatible with the principle of minimum relative entropy—a rather general principle of nature.
@@
We conclude by highlighting potential avenues for extending the findings of this paper. While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question. Establishing pure 2ESS’s asymptotic stability in multiplayer games using techniques beyond linear stability analysis is also missing. Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al. (2022) in non-matrix population games remains unexplored. Exploring the realm of asymmetric games in finite Nowak et al. (2004); Sekiguchi and Ohtsuki (2015) populations to locate an appropriate definition of ESS is also an interesting problem. Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al. (2023), opens up another avenue for investigation.
@@ Acknowledgements.
A subsequent sentence in the scheme of proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second. Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second. This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper. For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed NE), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$. Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$. To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if first component is zero then automatically second component is also zero. Let us see explicitly why this is so.
We know that any $2\times 2$ bimatrix game can be expressed in the form of either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V). For a rescaled zero-sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998). This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
In conclusion, if one accepts the theorem—present in Ref. Cressman (1996)—exactly as worded, then one is prone to erroneously deducing that mixed 2ESS in bimatrix games is asymptotically stable. This oversight in Ref. Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
@@ Appendix B ESS under Definition 2e for bihypermatrix games can only be pure
Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$. Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\delta}_{x},\mathbf{\hat{y}}+\bm{\delta}_{y})\equiv((\hat{x}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{y}+\delta_{y},1-\hat{y}-\delta_{y}))$. We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\delta_{x}$ and $\delta_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative. Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\delta}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\delta}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (88)
@@
$=$ $-\delta_{x}\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)-\delta_{y}\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
$=$ $-\delta_{x}\left(b_{1}\delta_{y}+b_{2}\delta_{y}^{2}+\cdots+b_{\gamma_{x}}\delta_{y}^{\gamma_{x}}\right)-\delta_{y}\left(c_{1}\delta_{x}+c_{2}\delta_{x}^{2}+\cdots+c_{\nu_{x}}\delta_{x}^{\nu_{y}}\right).$
@@
Now, to the leading order, the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\delta_{x}\delta_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\delta_{x}$ and $\delta_{y}$. For similar reason, in case $-(b_{1}+c_{1})\delta_{x}\delta_{y}=0$, the next order term in $\delta_{x}$ and $\delta_{y}$ of the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) leads to same conclusion, and so on. In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games. We remark the line of proof adopted herein can be extended to the case of mixed ESS under Definition 1e.